# From Benchmarks to Business Impact: Deploying IBM Generalist Agent in Enterprise Production

### v1

@@ 1 Introduction
  • Domain benchmark. The BPO-TA benchmark (26 tasks, 13 endpoints) that captures realistic enterprise analytics queries, enabling reproducible regression testing and ablation studies.
  • Architectural advances. A planner–executor agent design with schema-grounded prompting, variable tracking, reflective retries, provenance logging, and an API/Tool Hub that streamlined onboarding of enterprise applications. This architecture achieved state-of-the-art performance on both WebArena and AppWorld benchmarks, demonstrating its generality before adaptation for enterprise contexts.
@@ 2 Related Work
Early agentic paradigms such as ReAct interleave chain-of-thought reasoning with environment actions to improve task completion and interpretability (Yao et al. 2023), while code-centric approaches as CodeAct generate executable code to plan and call tools/APIs for complex tasks (Ye et al. 2024). These ideas catalyzed practical enterprise frameworks that orchestrate multiple specialized agents (or tools) with configurable roles. From AutoGen’s conversation-programmed multi-agent patterns (Wu et al. 2023), to LangGraph’s stateful, tool-grounded agent graphs for reliability and recoverability (LangChain 2024), and OpenAI’s Swarm orchestrator for multi-agent handoffs (OpenAI 2024). Despite promise, production experience consistently reports fragility at scale: brittle inter-agent handoffs, maintenance overhead from prompt/tool drift, safety and generalization across different domains.
@@
Enterprise report measurable wins when agentic systems are embedded behind assist and self-serve flows (The Verge 2025; Mobile World Live 2025; AI Business 2022). Analytics/BI copilots ship text-to-SQL/report-generation agents tightly coupled to enterprise data governance, observability, and review workflows (Snowflake Engineering 2024; Microsoft 2024; Databricks 2025). Broader surveys in hiring/HR analytics highlight fairness, transparency, and audit requirements in employment contexts, motivating provenance and explainability (Raghavan et al. 2020; Fabris et al. 2025; Chen 2023). Industry adoption reports likewise stress governance, measurable ROI, and operational readiness (monitoring, latency/cost budgets) as prerequisites for scale-out (Zhang et al. 2025b; Cemri et al. 2025). These constraints shape the design space of enterprise agents task completion. BPO process automation combine workflow orchestration, retrieval over enterprise knowledge, and document understanding to automate outsource non-core business functions to third-party providers.
@@
Concurrently, the research community has pushed toward generalist CUAs that plan and act across heterogeneous software. WebArena offers realistic, self-hosted websites for browser agents and showed early baselines struggled to exceed modest end-to-end success (Zhou et al. 2023). AppWorld evaluates multi-application orchestration via hundreds of programmatic APIs with outcome-based grading (Trivedi et al. 2024). OSWorld measures GUI workflows on desktop applications and OS tasks (Xie et al. 2024). Complementary suites probe interaction quality and oversight: ST-WebAgentBench emphasizes policy adherence in web agents, introducing Completion-under-Policy (CuP) as the primary objective (Levy et al. 2024), $\tau$-Bench targets tool-agent-user dynamics and policy/instruction following (Yao et al. 2024), BrowserGym provides a unified platform for evaluating web agents under controlled variability (de Chezelles et al. ). Architecturally, generalist CUAs increasingly combine hierarchical planning, explicit state/variable tracking, and reflective repair during execution to improve robustness in long-horizon settings (Shinn et al. 2023; Kim, Baldi, and McAleer 2023; Zhang et al. 2025c). Recent vendor-facing systems expose “computer use” capabilities (desktop/browser control, file ops) under sandboxes, signaling a trend toward production CUAs (Anthropic 2024; OpenAI 2025; Shen et al. 2025; Fourney et al. 2024; Marreed et al. 2025).
@@
Despite rapid progress, several gaps limit direct transfer from benchmark success to enterprise deployment. First, governance: high-risk domains (including employment analytics) demand provenance, HITL, and post-deployment monitoring mandated by frameworks and regulation (Tabassi 2023; European Union 2024; NYC Department of Consumer and Worker Protection 2023). Second, tool proliferation and schema variance: results degrade as agents shortlist from dozens of APIs and maintain consistency across dependent calls (Shen et al. 2024b; Xu et al. 2023). Third, operational constraints: latency, cost, and reproducibility must be tracked and controlled in production (Kwon et al. 2023; Zheng et al. 2024; Jiang et al. ). Fourth, graceful degradation: enterprise agents must decline unsupported requests without hallucination, and surface transparent rationales and computation logs (Nakano et al. 2021). Our work addresses this bridge by showing how a generalist CUA can be adapted for enterprise: schema-minimized API onboarding, deterministic parsing/validation, provenance-first responses, in a piloted BPO talent acquisition setting.
@@ 3.1 The Development Journey of Agentic Systems
The BPO–TA team followed this arc. An initial prototype, built quickly on reactive patterns, showed promise but could not scale to the breadth of sources and policies in Talent Acquisition. With 13 APIs spanning multiple systems and providers–each offering several actions and requiring orchestration across workflows–the complexity exceeded what the prototype could handle. At this inflection point, the team turned to IBM Research to evaluate the Computer Using Generalist Agent (CUGA), which had just achieved state-of-the-art results on AppWorld and WebArena benchmarks. The open question was: Could a benchmark-proven generalist agent deliver enterprise-grade performance in the demanding Talent Acquisition setting?
@@ 4 Enterprise Requirements for Generalist Agents
Enterprises will not adopt agents in production workflows unless they are confident the systems are safe and trustworthy, including:
@@
Enterprise processes are not generic. In Talent Acquisition, for example, the stages of requisition management and SLA definitions provide concrete guardrails. To succeed, agents must be able to align with such processes by:
  • Being taught via documents: ingesting enterprise playbooks, policies, and guidelines expressed in natural language.
  • Being shown via demonstrations: learning correct behavior from user examples.
  • Learning via feedback: adapting over time from implicit and explicit corrections, thereby refining their adherence to organizational norms.
@@ 5 System Architecture and Pre-deployment
CUGA implements a hierarchical agentic architecture with nested planner–executor loops (Fig. 2). At the top, an optional chat layer provides input interpretation and lightweight preprocessing, including message and variable histories; this can be bypassed in non-chat deployments. The outer loop governs task planning and orchestration: a Task Analyzer identifies the target application, a Task Decomposer determines whether multi-application coordination is required, and a persistent Plan Controller advances a durable task ledger. The ledger records steps, variable bindings, replans, and completions. In pilot evaluations, this ledger was essential for traceability, compliance, and recovery from partial failures. The inner loop delegates sub-tasks to specialized agents–API/Tool, Web Browser, CLI, and domain-specific agents–each acting within its own environment and returning structured observations to the controller.
@@
Two execution families illustrate the planner–executor pattern: the API Sub Agent and the Browser Sub Agent (Fig. 6, Fig. 5). The API Sub Agent combines short-term memory, an API Planner, and strategic reflection, coordinating a ShortlisterAgent (which selects APIs through a registry) and a CodeAgent (with a nested CodePlanner and sandboxed executor). This modular design allowed rapid onboarding of analytics endpoints in the BPO pilot. The Browser Sub Agent pairs a Browser Planner and Reflection Judge with two execution paths: an Action Agent (click, type, select, navigate) and a Question Answering Agent (DOM-to-Markdown conversion and screenshots). Although disabled in the BPO deployment for governance reasons, this design enables seamless switching between API-first and hybrid browser–API workflows without re-architecting.
@@ 6 Benchmarks and Pilot Evaluation
CUGA achieves state-of-the-art results on both WebArena and AppWorld, ranking first among published agents. Tables 1 and 2 present detailed per-application and per-level outcomes. On WebArena, CUGA attains an overall accuracy of 61.7%, with strongest performance on Reddit (75.5%) and Map (64.2%). On AppWorld’s ”Test-Challenge” dataset, CUGA reaches 48.2% overall scenario completion, with particularly high success on Level 1 tasks (87.5% scenario completion). These results demonstrate that the CUGA architecture is competitive with–and in many cases surpasses–specialized systems. More detilas can be found in Appendix C and D
@@
Table 2: Performance of CUGA on the AppWorld benchmark.
@@ 6.1 Enterprise Pilot: BPO-TA
While CUAs often perform well on abstract research benchmarks, enterprise deployment requires systematic offline evaluation with realistic tasks, audit guarantees, and compliance controls. To address this, we developed BPO-TA, a domain benchmark centered on talent acquisition (TA) workflows in the BPO context. BPO-TA encodes decision-support tasks drawn directly from analyst practice into a fixed test set, providing a reproducible regression baseline for measuring progress and ensuring operational reliability. Its design follows three principles essential for adoption: traceability (each task paired with APIs, glue code, and gold-standard explanations), realism (tasks grounded in genuine analyst workflows rather than synthetic probes), and reproducibility (fixed inputs, deterministic evaluation, and explicit provenance).
@@
Figure 3: Illustration of CUGA integrated into the BPO Talent Acquisition UI, providing natural-language analytics, provenance-aware explanations, and audit-ready compliance within enterprise workflows.
@@ 6.2 Potential Benefits
The CUGA system has been piloted within IBM’s Business-Process-Outsourcing (BPO) talent acquisition workflow since mid-2025. It can be used by recruiters and analyst teams to answer sourcing, funnel, and skill-impact questions that previously required manual data pulls and spreadsheet manipulation. The pilot was performed in a read-only configuration: CUGA connects to 13 domain-specific analytics APIs, each exposing pre-approved metrics such as service-level-agreement (SLA) performance, funnel conversions, and hires-by-source. Provenance logging and computation traces are stored for each interaction, ensuring audit readiness and compliance with organizational governance requirements (e.g., PII avoidance, immutable records of all API calls).
Preliminary evaluations of CUGA in simulated enterprise workflows suggest promising efficiency and reliability gains. Estimated benefits include a potential reduction in average time-to-answer (from roughly 20 minutes of manual work to an expected 2–5 minutes with CUGA, an estimated $\sim$90 % improvement) and higher reproducibility of responses (CUGA outputs were consistent across runs in about 90 % of internal test cases). Audit readiness is also expected to improve, with over 90 % of generated responses including full provenance (API endpoint, parameters, and result logs).
@@ 7 Lessons Learned and Insights
From this pilot, we derived a set of lessons and insights–both technical and organizational–that shape the path forward:
@@
8 Roadmap
The first phase surfaced clear requirements for the next stage of development. Key technical enhancements now underway include:
  • Configurable HITL. Fine-grained control over when agents act autonomously or defer to humans.
  • Safety and policy adherence. As the system expands from read-only analytics to autonomous actions, explicit mechanisms for enforcing enterprise policies and governance are critical.
  • Cost–performance tradeoffs. Cost and latency should be improved. Static and dynamic short-circuiting is being developed to reduce unnecessary reasoning by adapting the depth of reasoning to task complexity.
  • Trajectory reuse. Successful execution traces can be generalized into reusable tools, improving both reliability and efficiency when similar tasks recur.
  • Model optimization. Introducing smaller models for routine subtasks, while reserving large models for complex reasoning, balances accuracy with latency and cost.
A near-term milestone is to extend CUGA’s evaluation to TauBench2, which emphasizes policy adherence and human-in-the-loop operation. Success here will further validate CUGA as not only a benchmark-leading agent but also an enterprise-ready system that meets organizational standards of safety and trust.
@@ Appendix C WebArena Benchmark Analysis
The WebArena (Zhou et al. 2023) environment is a realistic and reproducible web environment designed for web agents to interact with. Along with the environment, a benchmark that consists of a set of 812 real-life natural language tasks has been released.
@@
Operator (OpenAI 2025) 58.1
Autonomous Web Agent (Jace.AI 2024) 57.1
ScribeAgent + GPT-4o (Shen et al. 2024a) 53
AgentSymbiotic (Zhang et al. 2025a) 52.1
@@ Appendix D AppWorld Dataset
AppWorld (Trivedi et al. 2024) is a high-quality execution environment designed to test agents with code generation capabilities. The AppWorld benchmark consists of 9 day-to-day application such as Spotify, Amazon Shopping, Gmail, available via 457 APIs and simulating the lives of simulated users. The benchmark is divided into two levels of difficulty - named ”Test-Normal” and ”Test-Challenge”, with the latter being generally more challenging, with each level consisting of several scenario data points, with each scenario consisting of several tasks for an agent to accomplish.
@@
Chen et al. (2025) Qwen2.5 72.6 (53.6) 47.2 (28.8)
Gupta et al. (2025) GPT-4o 68.5 (57.1) 38.9 (23)
@@ 26 Unsupported cross-app question (mixing HR + sourcing) None Must decline gracefully 0
Appendix F Experimental Results and Trajectories
This section presents detailed experimental results and example system trajectories demonstrating CUGA’s performance across different benchmarks.
Figure 10: Detailed performance results on the BPO-TA benchmark showing CUGA’s effectiveness across different task categories.
Figure 11: Detailed results on BPO benchmark - Part 1
Figure 12: CUGA performing the task: ”For each product type in my Amazon cart, buy the cheapest product and move the rest to the wish list.” The reflection API agent then summarized the progress so far and recommended moving non-cheapest products from the wish list as well, to avoid accidentally purchasing them later. This way, the agent doesn’t mistakenly buy products that are not the cheapest option.
Figure 13: Trajectory showing CUGA working in a hybrid mode that involves both the browser sub-agent and the API sub-agent. The task was: ’Get the top account by revenue from digital sales, then add it to the current page.’
Figure 14: The trajectory shows CUGA succeeded a task in the WebArena benchmark using the browser sub-agent.

### v2

@@ 1 Introduction
  • Domain benchmark. The BPO-TA benchmark that captures realistic enterprise analytics queries, enabling reproducible regression testing and ablation studies.
  • Architectural advances. A planner–executor agent design with schema-grounded prompting, variable tracking, reflective retries, provenance logging, and an API/Tool Hub that streamlined onboarding of enterprise applications. This architecture achieved state-of-the-art performance on both WebArena and AppWorld benchmarks.
@@ 2 Related Work
Early agentic paradigms such as ReAct interleave chain-of-thought reasoning with environment actions to improve task completion and interpretability (yao2023react), while code-centric approaches as CodeAct generate executable code to plan and call tools/APIs for complex tasks (ye2024codeact). These ideas catalyzed practical enterprise frameworks that orchestrate multiple specialized agents (or tools) with configurable roles. From AutoGen’s conversation-programmed multi-agent patterns (wu2023autogen), to LangGraph’s stateful, tool-grounded agent graphs for reliability and recoverability (langgraph2024), and OpenAI’s Swarm orchestrator for multi-agent handoffs (openaiSwarm2024). Despite promise, production experience consistently reports fragility at scale: brittle inter-agent handoffs, maintenance overhead from prompt/tool drift, safety and generalization across different domains.
@@
Enterprise report measurable wins when agentic systems are embedded behind assist and self-serve flows (verizonVerge2025; verizonMWL2025; aibusinessAlibaba2022). Analytics/BI copilots ship text-to-SQL/report-generation agents tightly coupled to enterprise data governance, observability, and review workflows (snowflakeCortex2024; msPowerBI2024; databricksGenie2025). Broader surveys in hiring/HR analytics highlight fairness, transparency, and audit requirements in employment contexts, motivating provenance and explainability (raghavan2020mitigating; fabris2025fairness; chen2023ethics; schwartz2023enhancing). Industry adoption reports likewise stress governance, measurable ROI, and operational readiness (monitoring, latency/cost budgets) as prerequisites for scale-out (zhang2025agent; cemri2025multi). These constraints shape the design space of enterprise agents task completion. BPO process automation combine workflow orchestration, retrieval over enterprise knowledge, and document understanding to automate outsource non-core business functions to third-party providers, leveraging semantic reasoning for process understanding as explored in (oved2025snap), and human-centric automation approaches such as IDA and conversational RPA (shlomov2024ida; yaeli2022recommending; zeltyn2022prescriptive).
@@
Concurrently, the research community has pushed toward generalist CUAs that plan and act across heterogeneous software. WebArena offers realistic, self-hosted websites for browser agents and showed early baselines struggled to exceed modest end-to-end success (zhou2023webarena). AppWorld evaluates multi-application orchestration via hundreds of programmatic APIs with outcome-based grading (trivedi2024appworld). OSWorld measures GUI workflows on desktop applications and OS tasks (xie2024osworld). Complementary suites probe interaction quality and oversight: ST-WebAgentBench emphasizes policy adherence in web agents, introducing Completion-under-Policy (CuP) as the primary objective (levy2025stwebagentbench), $\tau$-Bench targets tool-agent-user dynamics and policy/instruction following (yao2024tau), BrowserGym provides a unified platform for evaluating web agents under controlled variability (debrowsergym) while (shlomov2024grounding) identifies planning as the dominant bottleneck in web agents. Architecturally, generalist CUAs increasingly combine hierarchical planning, explicit state/variable tracking, and reflective repair during execution to improve robustness in long-horizon settings (shinn2023reflexion; kim2023language; zhang2025agentorchestra). Recent vendor-facing systems expose “computer use” capabilities (desktop/browser control, file ops) under sandboxes, signaling a trend toward production CUAs (anthropic2024computeruse; openai2025operator; shen2025mind; fourney2024magentic).
@@
Despite rapid progress, several gaps limit direct transfer from benchmark success to enterprise deployment. First, governance: high-risk domains (including employment analytics) demand provenance, HITL, and post-deployment monitoring mandated by frameworks and regulation (nist2023airmf; euai2024; nyc144). Second, tool proliferation and schema variance: results degrade as agents shortlist from dozens of APIs and maintain consistency across dependent calls (shen2024taskbench; xu2023tool). Third, operational constraints: latency, cost, and reproducibility must be tracked and controlled in production (kwon2023efficient; zheng2024sglang; jiang2023llmlingua). Fourth, graceful degradation: enterprise agents must decline unsupported requests without hallucination, and surface transparent rationales and computation logs (nakano2021webgpt). Our work addresses this bridge by showing how a generalist CUA can be adapted for enterprise: schema-minimized API onboarding, deterministic parsing/validation, provenance-first responses, in a piloted BPO talent acquisition setting.
@@ 3.1 The Development Journey of Agentic Systems
The BPO–TA team followed this arc. An initial prototype, built quickly on reactive patterns, showed promise but could not scale to the breadth of sources and policies in Talent Acquisition. With 13 APIs spanning multiple systems and providers–each offering several actions and requiring orchestration across workflows–the complexity exceeded what the prototype could handle. At this inflection point, the team turned to IBM Research to evaluate CUGA, which had just achieved state-of-the-art results on AppWorld and WebArena benchmarks. The open question was: Could a benchmark-proven generalist agent deliver enterprise-grade performance in the demanding Talent Acquisition setting?
@@ 4 Enterprise Requirements for Generalist Agents
Enterprises adopt agents in production only when their safety and trustworthiness are ensured, including:
@@
Enterprise processes are rarely generic. In Talent Acquisition, for example, requisition workflows and SLA definitions act as concrete guardrails that agents must follow. Effective alignment requires agents to learn from enterprise documentation (such as playbooks, policies, and guidelines written in natural language), from user demonstrations that model correct behavior, and from ongoing feedback that enables them to gradually refine their responses and conform to organizational norms.
@@ 5 System Architecture and Pre-deployment
CUGA implements a hierarchical agentic architecture (marreed2025towards) with nested planner–executor loops (Fig. 2). At the top, an optional chat layer provides input interpretation and lightweight preprocessing, including message and variable histories; this can be bypassed in non-chat deployments. The outer loop governs task planning and orchestration: a Task Analyzer identifies the target application, a Task Decomposer determines whether multi-application coordination is required, and a persistent Plan Controller advances a durable task ledger. The ledger records steps, variable bindings, replans, and completions. In pilot evaluations, this ledger was essential for traceability, compliance, and recovery from partial failures. The inner loop delegates sub-tasks to specialized agents–API/Tool, Web Browser, CLI, and domain-specific agents–each acting within its own environment and returning structured observations to the controller.
@@
Two execution families illustrate the planner–executor pattern: the API Sub Agent and the Browser Sub Agent. The API Sub Agent combines short-term memory, an API Planner, and strategic reflection, coordinating a ShortlisterAgent (which selects APIs through a registry) and a CodeAgent (with a nested CodePlanner and sandboxed executor). This modular design allowed rapid onboarding of analytics endpoints in the BPO pilot. The Browser Sub Agent pairs a Browser Planner and Reflection Judge with two execution paths: an Action Agent (click, type, select, navigate) and a Question Answering Agent (DOM-to-Markdown conversion and screenshots). Although disabled in the BPO deployment for governance reasons, this design enables seamless switching between API-first and hybrid browser–API workflows without re-architecting.
@@ 6 Benchmarks and Pilot Evaluation
CUGA achieves state-of-the-art results on both WebArena and AppWorld, ranking first among published agents. Tables 1 and 2 present detailed per-application and per-level outcomes. On WebArena, CUGA attains an overall accuracy of 61.7%, with the strongest performance on Reddit (75.5%) and Map (64.2%). On AppWorld’s ”Test-Challenge” dataset, CUGA reaches 48.2% overall scenario completion, with particularly high success on Level 1 tasks (87.5% scenario completion). These results demonstrate that the CUGA architecture is competitive with–and in many cases surpasses–specialized systems. More details can be found in the Appendix.
@@
Table 2: Performance of CUGA on AppWorld.
@@ 6.1 Enterprise Pilot: BPO-TA
While CUGAs often perform well on abstract research benchmarks, enterprise deployment requires systematic offline evaluation with realistic tasks, audit guarantees, and compliance controls. To address this, we developed BPO-TA, a domain benchmark centered on talent acquisition (TA) workflows in the BPO context. BPO-TA encodes decision-support tasks drawn directly from analyst practice into a fixed test set, providing a reproducible regression baseline for measuring progress and ensuring operational reliability. Its design follows three principles essential for adoption: traceability (each task paired with APIs, glue code, and gold-standard explanations), realism (tasks grounded in genuine analyst workflows rather than synthetic probes), and reproducibility (fixed inputs, deterministic evaluation, and explicit provenance).
@@ 6.2 Potential Benefits
The CUGA system has been piloted within IBM’s Business-Process-Outsourcing (BPO) talent acquisition workflow since mid-2025. It can be used by recruiters and analyst teams to answer sourcing, funnel, and skill-impact questions that previously required manual data pulls and spreadsheet manipulation. The pilot was performed in a read-only configuration: CUGA connects to 13 domain-specific analytics APIs, each exposing pre-approved metrics such as funnel conversions and hires-by-source. Provenance logging and computation traces are stored for each interaction, ensuring audit readiness and compliance with organizational governance requirements (e.g., PII avoidance, immutable records of all API calls).
Preliminary evaluations of CUGA in simulated enterprise workflows, although not formally tested for statistical significance (dror2018hitchhiker; dror2020statistical), suggest promising efficiency and reliability gains. suggest promising efficiency and reliability gains. Estimated benefits include a potential reduction in average time-to-answer (from roughly 20 minutes of manual work to an expected 2–5 minutes with CUGA, an estimated $\sim$90 % improvement) and higher reproducibility of responses (CUGA outputs were consistent across runs in about 90 % of internal test cases). Audit readiness is also expected to improve, with over 90 % of generated responses including full provenance (API endpoint, parameters, and result logs).
@@ 7 Lessons Learned and Insights
From this pilot, we derived a set of technical and organizational lessons that shape the path forward:
@@
CUGA’s first phase surfaced clear requirements for enterprise deployment, driving current work on configurable human-in-the-loop control, explicit policy-enforcement for safe autonomous actions, improved cost–latency tradeoffs through adaptive short-circuiting, reuse of successful trajectories as tools, and selective use of smaller models for routine tasks. The next milestone is evaluation on policy compliance and HITL governance as an enterprise-ready system that meets organizational standards of safety and trust.
@@ Appendix C WebArena Benchmark Analysis
The WebArena (zhou2023webarena) environment is a realistic and reproducible web environment designed for web agents to interact with. Along with the environment, a benchmark that consists of a set of 812 real-life natural language tasks has been released.
@@
Operator (openai2025operator) 58.1
Autonomous Web Agent (jaceAIWebArenaResults) 57.1
ScribeAgent + GPT-4o (shen2024scribeagent) 53
AgentSymbiotic (zhang2025symbiotic) 52.1
@@ Appendix D AppWorld Dataset
AppWorld (trivedi2024appworld) is a high-quality execution environment designed to test agents with code generation capabilities. The AppWorld benchmark consists of 9 day-to-day application such as Spotify, Amazon Shopping, Gmail, available via 457 APIs and simulating the lives of simulated users. The benchmark is divided into two levels of difficulty - named ”Test-Normal” and ”Test-Challenge”, with the latter being generally more challenging, with each level consisting of several scenario data points, with each scenario consisting of several tasks for an agent to accomplish.
@@
chen2025reinforcement Qwen2.5 72.6 (53.6) 47.2 (28.8)
gupta2025leveraging GPT-4o 68.5 (57.1) 38.9 (23)